Task list for all observatories

20 Inch F4 in Dome task list

  • ACP plans need refining to remove objects that are not galaxies or are too small/faint (ONGOING 80% complete as of Nov 2010).
  • Dome can stick with ice to the skirt in very cold weather. Old the skirt to the dome rub ring. When this happens you get bad scope comms in he Sitech controller as the scope has run into the dome az limit and cut the power. You need to move the dome physically back in sync with the scope. PROBLEM SOLVED. Grease the skirt to the dome wall. White grease sprayed on. PROBELM REAPPEARED. The dome skirt has now been added in sections so it is not too tight. COMPLETED
  • Dome Azimuth motor too slow. Need to replace to increase slew speed to >2 deg/sec. VSD required as in run off roof to replace shunt motor.
  • Absolute encoders to be installed
  • Motorised Focus(now under control using the Observatory Auxiliary Control sofwtare. COMPLETED)
  • Derotator drive Sitech needed (not essential at the moment with short exposures for Supernova runs)
  • ACP to Sitech offset VERY important! Still not fixed. Dan has been emailed. No response as yet. PROBLEM SOLVED. New version V0.53H has fixed the problem.
  • Indictaor required to state doem open or closed and/or ok to open the dome. Web cam be able to see indicator. (COMPLETED, new Observatory Auxiliary control softare and internal web cam)
  • Dome door drive too slow. New more powerful and faster drive required. (this has now been upgraded with a new motor. COMPLETED)
  • Plan chaining in ACP. Dry run first in the day with any plate solves. Not now required. New larger plans being developed
  • Add SHUTDOWN to the ACP plans (this is added to the last or approprite plan (SEE USER GUIDE)
  • Replace the main rain sensor and add additional for backup. Add shroudd to protect from direct sunlight.
  • ASCOM problem with ACP/Sitech.? Dan/Bob do not know what the problem is. Fix at the moment is to bin out Sitech immediately after starting ACP run (TIMER EXCEPTION) then restart Sitech and reconnect to scope in ACP. roblem fixed. Not great!
  • ACP Script added to root of ACP to #SHUTDOWN if a plan fails COMPLETED
  • Replace skirt and fixing DONE
  • Clean and wash dome DONE
  • Water seal brickwork outside DONE
  • Clean inside and paint red DONE
  • Dew protector for notebook to replace temp cover Still outstanding
  • Tweak Alt drive system and align Needs reqork as still get altitude issues when tracking 20% images.

Wide field camera task list

  • Connect cable in control room to weather station 24V loop DONE
  • Pull new green UTP Cat 5 cable from control room to observatory and terminate bothe nds then test DONE
  • Check IP of machine and ensure that the router is configured for port forwarding DONE but needs review with Zej
  • Reload ACP and GSC and configure DONE
  • Check park position and set to custom position that is pointing to the pole DONE
  • Additional fixings to side panels DONE
  • Sound proof box and finger protection guard for motor DONE
  • Fix main structure to floor DONE
  • Fit focal extender and DONE no reducer installed now new CCD Atik 320e is installed with new focuser
  • L strip and double P seal to bottom edge inside of walls DONE
  • Gutter down pipe NOT YET DONE!!!
  • Tidy cableing DONE
  • Enable dew heater and test temperature Dew heate now switched on. check temp over time
  • Move open limit and ultimate limit block to prevent cable damage and ensure can still see pole DONE
  • Build ACP plans for wide field camera Not yet done!!!
  • Dew protector for notebook DONE
  • Update documents and drawings and load up to wiki Only 70% done. drawings to update
  • THE ROOF RAILS ICE UP VERY EASILY AND PREVENT THE ROOF CLOSING. MAJOR ISSUE TO RESOLVE.